Suppose you have a sample of phosphorus-32 that has only 3.1 percent of its original radioactivity, how old is this sample if its half-life is fourteen days?
N/N0=.031

Answer by stanbon(75887)

Suppose you have a sample of phosphorus-32 that has only 3.1 percent of its original radioactivity, how old is this sample if its half-life is fourteen days?
---
0.031No = No(1/2)^(t/14)
---
0.031 = (1/2)^(t/14)
---
(t/14) = log(0.031)/log(1/2)
t/14 = 5.0116
time = 70.1622 days
==========================
Cheers,
Stan H.